# On an entropy inequality for quadratic forms and applications

### v1

@@ 3 Discretized entropy Balog-Szemerédi-Gowers theorem
In this section, we follow the method introduced by Tao in [22, Theorem 3.1] to obtain a generalized discretized entropy version of the Balog-Szemerédi-Gowers theorem. We do not know whether such a generalization has been studied in the literature, so we provide a complete proof in this section. The Balog-Szemerédi-Gowers theorem (or one of its many variants) says that if one has a finite subset of the integers whose sums line up a lot, then it must not have many sums. More precisely, given $A\subseteq\mathbb{Z},$ whose elements $a_{j}\in A$ have $\gamma|A|^{3}$ relations of the form $a_{1}+a_{2}=a_{3}+a_{4},$ there must be a large subset $A^{\prime}\subseteq A,$ where $|A^{\prime}|\geq C_{1}(\gamma)|A|$ that satisfies $|A^{\prime}+A^{\prime}|\leq C_{2}(\gamma)|A^{\prime}|,$ where the constants $C_{j}$ above are positive and only depend on $\gamma.$
@@
The original version of entropy Balog-Szemerédi-Gowers theorem in [22] states for $H$ instead of $H_{n}$, does not involve the functions $f$, $g$, and conditions (ii)–(iv) are “given $Y^{\prime}$, $X^{\prime}$, $X^{\prime}$” instead of “given $D_{n}(Y^{\prime})$, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime})$”, respectively. In such a situation, (i)–(iv) is equivalent to the statement that $X^{\prime},Y,X^{\prime\prime},Y^{\prime\prime}$ forms a Markov chain.
@@
Applying Lemma B.1 with $A=0$, $B=X$, $C=Y$ and $h=D_{n}$, we obtain a random variable $(X_{1},Y_{1},X_{2})$ such that: $(X_{1},Y_{1})\equiv(X,Y)$; $X_{2}$ and $X_{1}$ are conditionally i.i.d. given $D_{n}(Y_{1})$.
Next, let us apply Lemma B.1 again, with $A=X_{2}$, $B=Y_{1}$, $C=X_{1}$ and $h=D_{n}$. There exists a random variable $(X_{2}^{\prime},Y_{1}^{\prime},X_{1}^{\prime},Y_{1}^{\prime\prime})$ such that: $(X_{2}^{\prime},Y_{1}^{\prime},X_{1}^{\prime})\equiv(X_{2},Y_{1},X_{1})$, $(Y_{1}^{\prime\prime},D_{n}(X_{1}^{\prime}))\equiv(Y_{1},D_{n}(X_{1}))$; $Y_{1}^{\prime\prime}$ are $Y_{1}^{\prime}$ are conditionally i.i.d. given $D_{n}(X_{1}^{\prime})$; also $(X_{2}^{\prime},Y_{1}^{\prime})$ and $Y_{1}^{\prime\prime}$ are conditionally independent given $D_{n}(X_{1}^{\prime})$.
One can also use $A=(X_{2},X_{1},D_{n}(Y_{1}))$ in the second application of Lemma B.1 to see $(X_{2}^{\prime},X_{1}^{\prime},D_{n}(Y_{1}^{\prime}))\equiv(X_{2},X_{1},D_{n}(Y_{1}))$. Now (i)–(iv) hold by taking $X^{\prime}=X_{1}^{\prime}$, $X^{\prime\prime}=X_{2}^{\prime}$, $Y^{\prime}=Y_{1}^{\prime}$, $Y^{\prime\prime}=Y_{1}^{\prime\prime}$.
@@
Note that $D_{n}(Z)$, $D_{n}(Y^{\prime})$ are conditionally i.i.d. given $D_{n}(X,X^{\prime})$. Also, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime\prime})$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$. By Lemmas 2.8 and 2.19,
@@
$\begin{split}&H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|Y^{\prime})=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})-H_{n}(Y^{\prime}).\\ \end{split}$ (26)
Using Lemmas 2.16 and 2.18, we have that
@@
4 Proof of Theorem 1.1 with $s>1/2$: diagonal forms
@@
Let $\phi(x,y)=\alpha x^{2}+\beta y^{2}$, where $\alpha,\beta\in\mathbb{R}^{\ast}$. Let $X$, $Y$ be independent random variables with $range_{+}(X)\cup range_{+}(Y)\subseteq[-c_{1},-c_{2}]\cup[c_{2},c_{1}]$, where $0<c_{2}<c_{1}$. Suppose that $X$ and $Y$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, with $11/2<s<$ 1 and $C_{1},C_{2}\geq 1$. Then there exists a constant $\epsilon_{1}=\epsilon_{1}(s)>0$ such that
$\max\{H_{n}(X+Y),~{}H_{n}(\phi(X,Y))\}\geq n(s+\epsilon_{1})+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)$
@@
Remark 4.2.
Compared to Theorem 1.1, this provides a stronger statement, namely, for all non-zero real-coefficients $\alpha$ and $\beta$.
@@
Write $\phi(x,y)=p(x)+q(y)$, where $p(x)=\alpha x^{2}$ and $q(y)=\beta y^{2}$ are functions whose domains are both taken to be $[-c_{2},-c_{1}]\cup[c_{2},c_{1}]$. One asserts by Example 2.17 that
$H_{n}(p(X))=H_{n}(X)+O_{\phi,c_{1},c_{2}}(1),\quad H_{n}(q(Y))=H_{n}(Y)+O_{\phi,c_{1},c_{2}}(1).$ (31)
@@
From Example 2.12 and the fact that $range(X_{1}+X_{2}),range(Y_{1}+Y_{2})\subseteq[-2c_{1},2c_{1}]$, one sees that $f_{1}$ is $O_{\phi,c_{1}}(1)$-Lipschitz. We also have that $|W_{1}-W_{2}|^{2}=f_{2}(X_{1}-X_{2},Y_{1}-Y_{2})$, with the $O_{\phi,c_{1}}(1)$-Lipschitz function
@@
$\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O_{\phi,c_{1}}(1).$ (32)
We now estimate the terms in the last expression. For the first term, apply Corollary 2.22 to get that
@@
For the third term, we use Corollary 2.22 repeatedly, and combine Lemma 2.16 and formula (31), to get
@@
$=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(Y)+O_{\phi,c_{1},c_{2}}(1).$ (34)
For the fourth term, applying Corollary 2.22 shows that
@@
$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))+O_{\phi,c_{1},c_{2}}(1).$
Note that $X$ and $Y$ are independent, and are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively. One asserts that $W$ is $(2s,C_{1}C_{2}(\alpha\beta)^{-s/2})$-Frostman. Noting that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O_{\phi,c_{1}}(1)$. So we can deduce by Lemma 2.26 and Theorem 2.35 that
$H_{n}\big{(}X\big{)}\geq ns-\log C_{1}+O(1),\quad H_{n}\big{(}Y\big{)}\geq ns-\log C_{2}+O(1),$
$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}\geq n(s+\epsilon_{0})+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)$
@@
$n\left(11s+\epsilon_{0}\right)+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y)).$
As a consequence,
$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}$ $\geq n\left(s+\frac{\epsilon_{0}}{11}\right)+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1).$
@@
$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+8H_{n}(X)+8H_{n}(Y)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))+O_{\phi,c_{1},c_{2}}(1),$
and
$n\left(17s+\epsilon_{0}\right)+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y)).$
@@
We emphasize that the parameters $C_{1},C_{2}$ in Lemma 4.1 can be unbounded as $n$ tends to infinity. And, when we apply the above lemma in the next section, the term $O(\log C_{1}C_{2})$ will lead to some entropy-terms. So it cannot be replaced by $O_{C_{1},C_{2}}(1)$.
5 Proof of Theorem 1.1 with $s>1/2$: general forms
In this section, we first consider some special quadratic forms that are more complicated than the diagonal ones. In the following lemma, when we say a constant depends on the quadratic form $\phi$, this means that it depends on the coefficients $\alpha$ and $\beta$ occurred in $\phi$.
Lemma 5.1.
Let $\phi(x,y)=\alpha(x-py)^{2}+\beta(x+y)^{2}$, where $\alpha,\beta\in{\mathbb{R}}^{\ast}$ and $p\in\{0,1\}$. Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $[-c_{1},c_{1}]$, where $11/2<s<$ 1, $C\geq 1$ and $c_{1}>0$. Then there exists a positive constant $\epsilon_{2}=\epsilon_{2}(s)$ and an $N=N(\phi,s,C,c_{1})$ such that
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{2})$
for all $n>N$.
@@
$\Psi:=\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}-ns,$ (36)
and our aim is to prove that $\Psi\geq\epsilon_{2}n$ for some positive $\epsilon_{2}$.
Since $X$, $Y$ are independent and both $(s,C)$-Frostman, we deduce by Lemma 2.26 Corollary 2.22 and (36) that
$\begin{split}&\min\{H_{n}(X),\,H_{n}(Y)\}\geq ns+O_{C}(1),\\ &\max\{H_{n}(X),\,H_{n}(Y)\}\leq H_{n}(X+Y)+O(1)\leq ns+\Psi+O(1).\end{split}$ (37)
Let us define the random variable $(X_{0},Y_{0})$ as
$(X_{0},Y_{0})=(X-pY,X+Y)$
with $p\in\{0,1\}$. The ranges of $X_{0}$ and $Y_{0}$ are both contained in $[-2c_{1},2c_{1}]$. And $(X_{0},Y_{0})$ is an $O(1)$-bi-Lipschitz function of $(X,Y)$, which, because $X$ and $Y$ are independent, leads to the estimate
$H_{n}(X_{0},Y_{0})=H_{n}(X,Y)+O(1)=H_{n}(X)+H_{n}(Y)+O(1)\leq 2ns+2\Psi+O(1).$ (38)
By Lemma 2.25, one deduces that $X_{0}$ and $Y_{0}$ are $(s,C_{1}^{\prime})$-Frostman and $(s,C_{2}^{\prime})$-Frostman, respectively, with $C_{1}^{\prime},C_{2}^{\prime}=O_{C}(1)$. Similarly, one also has that $X_{0}+Y_{0}=(2X+(1-p)Y)$ is $(s,O_{C}(1))$-Frostman, and $(X_{0},Y_{0})$ is $(2s,O_{C}(1))$-Frostman. We then show below that
$H_{n}(X_{0})\leq ns+O(\Psi)+O_{C}(1),\quad H_{n}(X_{0}+Y_{0})\leq ns+O(\Psi)+O_{C}(1).$ (39)
Indeed, when $p=0$, one sees by (36), (37) and Lemma 2.23 that
$H_{n}(X_{0})=H_{n}(X)\leq ns+\Psi+O(1),$
and
$H_{n}(X_{0}+Y_{0})=H_{n}(2X+Y)$
$\qquad\leq H_{n}(X+Y)+O(H_{n}(X+Y)-H_{n}(X))+O(H(X+Y)-H(Y))+O(1)$
$\qquad\leq ns+O(\Psi)+O_{C}(1).$
When $p=1$, it follows from (12), (36) and (37) that
$H_{n}(X_{0})=H_{n}(X-Y)\leq 3H_{n}(X+Y)-H(X)-H(Y)+O(1)\leq ns+3\Psi+O_{C}(1)$
and
$H_{n}(X_{0}+Y_{0})=H_{n}(2X)=H_{n}(X)+O(1)\leq ns+\Psi+O(1).$
So (39) follows.
[…]
$P(E)=P(|X_{0}|>\eta,|Y_{0}|>\eta)\geq 1-P(|X_{0}|<\eta)-P(|Y_{0}|<\eta)\geq 1-C_{1}^{\prime}\eta^{s}-C_{2}^{\prime}\eta^{s}>\frac{1}{2},$
Lemma B.2 shows that $(X_{1},Y_{1})$ is $(2s,O_{C}(1))$-Frostman. Similarly, all of $X_{1}$, $Y_{1}$ and $X_{1}+Y_{1}$ are $(s,O_{C}(1))$-Frostman. We can apply Lemmas B.2 and 2.26, then combine (36), (38), (39) to get that
$\begin{split}&2ns+O_{C}(1)\leq H_{n}(X_{1},Y_{1})\leq\frac{H_{n}(X_{0},Y_{0})}{P(E)}\leq(1+O_{C}(\eta^{s}))(2ns+2\Psi+O(1)),\\ &ns+O_{C}(1)\leq H_{n}(X_{1})\leq\frac{H_{n}(X_{0})}{P(E)}\leq(1+O_{C}(\eta^{s}))(ns+O(\Psi)+O_{C}(1)),\\ &ns+O_{C}(1)\leq H_{n}(Y_{1})\leq\frac{H_{n}(X+Y)}{P(E)}\leq(1+O_{C}(\eta^{s}))(ns+\Psi),\\ &ns+O_{C}(1)\leq H_{n}(X_{1}+Y_{1})\leq\frac{H_{n}(X_{0}+Y_{0})}{P(E)}\leq(1+O_{C}(\eta^{s}))(ns+O(\Psi)+O_{C}(1)),\end{split}$ (40)
and
$H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})\leq\frac{H_{n}(\alpha X_{0}^{2}+\beta Y_{0}^{2})}{P(E)}=(1+O_{C}(\eta^{s}))H_{n}(\phi(X,Y))\leq(1+O_{C}(\eta^{s}))(ns+\Psi).$ (41)
@@
$\begin{split}&H_{n}(X_{1})+H_{n}(Y_{1})-H_{n}(X_{1},Y_{1})=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1),\\ &H_{n}(X_{1}+Y_{1})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1),\\ &H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1).\end{split}$
Now we write $\log K:=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1)$. Applying Theorem 3.1 with the random variable $(X_{1},Y_{1})$, we obtain the random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$, where $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$, and
$H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(X_{1})-\log K,\quad H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(Y_{1})-\log K.\$ (42)

### v2

@@ 3 Discretized entropy Balog-Szemerédi-Gowers theorem
In this section, we follow the method introduced by Tao in [26, Theorem 3.1] to obtain a generalized discretized entropy version of the Balog-Szemerédi-Gowers theorem. We do not know whether such a generalization has been studied in the literature, so we provide a complete proof in this section. The Balog-Szemerédi-Gowers theorem (or one of its many variants) says that if one has a finite subset of the integers whose sums line up a lot, then it must not have many sums. More precisely, given $A\subseteq\mathbb{Z},$ whose elements $a_{j}\in A$ have $\gamma|A|^{3}$ relations of the form $a_{1}+a_{2}=a_{3}+a_{4},$ there must be a large subset $A^{\prime}\subseteq A,$ where $|A^{\prime}|\geq C_{1}(\gamma)|A|$ that satisfies $|A^{\prime}+A^{\prime}|\leq C_{2}(\gamma)|A^{\prime}|,$ where the constants $C_{j}$ above are positive and only depend on $\gamma.$
@@
The original version of entropy Balog-Szemerédi-Gowers theorem in [26] states for $H$ instead of $H_{n}$, does not involve the functions $f$, $g$, and conditions (ii)–(iv) are “given $Y^{\prime}$, $^{\prime}, XX^{\prime}$” instead of “given $D_{n}(Y^{\prime})$, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime})$”, respectively. In such a situation, (i)–(iv) is equivalent to the statement that $X^{\prime},Y,X^{\prime\prime},Y^{\prime\prime}$ forms a Markov chain.
@@
Applying Lemma A.1 with $A=0$, $B=X$, $C=Y$ and $h=D_{n}$, we obtain a random variable $(X_{1},Y_{1},X_{2})$ such that: $(X_{1},Y_{1})\equiv(X,Y)$; $X_{2}$ and $X_{1}$ are conditionally i.i.d. given $D_{n}(Y_{1})$.
Next, let us apply Lemma A.1 again, with $A=X_{2}$, $B=Y_{1}$, $C=X_{1}$ and $h=D_{n}$. There exists a random variable $(X_{2}^{\prime},Y_{1}^{\prime},X_{1}^{\prime},Y_{1}^{\prime\prime})$ such that: $(X_{2}^{\prime},Y_{1}^{\prime},X_{1}^{\prime})\equiv(X_{2},Y_{1},X_{1})$, $(Y_{1}^{\prime\prime},D_{n}(X_{1}^{\prime}))\equiv(Y_{1},D_{n}(X_{1}))$; $Y_{1}^{\prime\prime}$ are $Y_{1}^{\prime}$ are conditionally i.i.d. given $D_{n}(X_{1}^{\prime})$; also $(X_{2}^{\prime},Y_{1}^{\prime})$ and $Y_{1}^{\prime\prime}$ are conditionally independent given $D_{n}(X_{1}^{\prime})$.
One can also use $A=(X_{2},X_{1},D_{n}(Y_{1}))$ in the second application of Lemma A.1 to see $(X_{2}^{\prime},X_{1}^{\prime},D_{n}(Y_{1}^{\prime}))\equiv(X_{2},X_{1},D_{n}(Y_{1}))$. Now (i)–(iv) hold by taking $^{\prime}=X_{1}^{\prime}, X′′=X$$^{\prime\prime}=X_{2}^{\prime}$, $^{\prime}=Y_{1}^{\prime}, Y′′=Y$$^{\prime\prime}=Y_{1}^{\prime\prime}$.
@@
Note that $D_{n}(Z)$, $D_{n}(Y^{\prime})$ are conditionally i.i.d. given $D_{n}(X,X^{\prime})$. Also, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime\prime})$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$. By Lemmas 2.8 and 2.18,
@@
[…]
Using Lemmas 2.15 and 2.17, we have that
@@
4 Proof of Theorems 1.2-1.4 with $s>1/2$
4.1 Diagonal forms with independent random variables
@@
Let $\alpha,\beta\in\mathbb{R}^{\ast}$, $1/2<s<1$, $C_{1},C_{2}>0$ and $0<c_{2}<c_{1}$. There exists a constant $\epsilon_{1}=\epsilon_{1}(s)>0$ such that the following statement holds: Let $X$, $Y$ be independent random variables with $range_{+}(X)\cup range_{+}(Y)\subseteq[-c_{1},-c_{2}]\cup[c_{2},c_{1}]$. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\begin{split}&\max\{H_{n}(X+Y),~H_{n}(\alpha X^{2}+\beta Y^{2})\}\geq n(s+\epsilon_{1})\\ &\qquad\qquad\qquad+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).\end{split}$ (35)
@@
Write $\phi(x,y)=p(x)+q(y)$, where $p(x)=\alpha x^{2}$ and $q(y)=\beta y^{2}$ are functions whose domains are both taken to be $c_{2},-c_{1}]\cup[c_{2},c_{1}]$. One asserts by Example 2.16 that
$\begin{split}H_{n}(p(X))&=H_{n}(X)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1),\\ H_{n}(q(Y))&=H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).\end{split}$ (36)
@@
From Example 2.11 and the fact that $range(X_{1}+X_{2}),range(Y_{1}+Y_{2})\subseteq[-2c_{1},2c_{1}]$, one sees that $f_{1}$ is $O_{\phi}(c_{1})$-Lipschitz. We also have that $|W_{1}-W_{2}|^{2}=f_{2}(X_{1}-X_{2},Y_{1}-Y_{2})$, with the $O_{\phi}(C_{1})$-Lipschitz function
@@
$\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O(\log_{+}c_{1})+O_{\phi}(1).$ (37)
We now estimate the terms in the last expression. For the first term, apply Lemma 2.20 to get that
@@
For the third term, we use Lemma 2.20 repeatedly, and combine Lemma 2.15 and formula (36), to get
@@
$=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$ (39)
For the fourth term, applying Lemma 2.20 shows that
@@
$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+$ $5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))$
$\quad+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$
Since $X$ and $Y$ are independent, and are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively, one asserts that $W$ is $(s_{1}+s_{2},C_{1}C_{2}\alpha^{-s_{1}}\beta^{-s_{2}})$-Frostman. Here $\log(\alpha^{-s_{1}}\beta^{-s_{2}})=O_{\phi}(1)$. Noting that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O_{\phi,c_{1}}(1)$. So we can deduce by Lemma 2.24 and Theorem 2.33 that
$H_{n}\big{(}X\big{)}\geq ns_{1}-\log C_{1}+O(1),\quad H_{n}\big{(}Y\big{)}\geq ns_{2}-\log C_{2}+O(1),$
$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}\geq n(s+\epsilon_{0})+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O_{\phi,s}(1)$
@@
$n\left(s+\epsilon_{0}+5s_{1}+5s_{2}\right)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))$
$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).$
Recalling that $s_{1}+s_{2}\geq 2s$, we conclude that
$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n\left(s+\frac{\epsilon_{0}}{11}\right)$
$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).$
@@
$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+$ $8H_{n}(X)+8H_{n}(Y)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))$
$\quad+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1),$
and then
$n\left(s+\epsilon_{0}+8s_{1}+8s_{2}\right)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))$
$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).$
@@
We emphasize that the parameters $C_{1},C_{2}$ in Lemma 4.1 can be unbounded as $n$ tends to infinity. And, when we apply the above lemma in the next subsection, the term $O(\log C_{1}C_{2})$ will lead to some entropy-terms. So it cannot be replaced by $O_{C_{1},C_{2}}(1)$.
4.2 Diagonal forms with dependent random variables
Now we prove Theorem 1.4 in the following form, where $X$ and $Y$ are allowed to be dependent. We still take $\phi(x,y)=\alpha x^{2}+\beta y^{2}$ in this subsection. And a constant depending on $\phi$ means it depends on $\alpha$ and $\beta$.
Theorem 1.4’. Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$. There is a positive constant $\epsilon_{2}=\epsilon_{2}(s)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1).$
@@ 4.2 Diagonal forms with dependent random variables
$\Psi:=\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}-ns,$
and our aim is to prove that $\Psi\geq\epsilon n$ for some positive $\epsilon$. In particular,
$\frac{H_{n}(X)+H_{n}(Y)}{2},\,\,H_{n}(X+Y),\,\,H_{n}(\alpha X^{2}+\beta Y^{2})\,\leq\,ns+\Psi.$ (41)
By Lemma 2.22, the random variables $X$, $Y$, $(X,Y)$ are $(s_{1},C\cdot\max\{1,c\})$-Frostman, $(s_{2},C\cdot\max\{1,c\})$-Frostman and $(s_{1}+s_{2},C)$-Frostman, respectively.
Let $\eta$ be a parameter to be chosen later, with $0<\eta<\min\{2^{-1}(1/2C)^{1/2s},\,1/10\}$. We denote by $E$ the event that $|(X,Y)|\geq\eta$, and by $P_{E}$ the corresponding conditional probability measure. Denote by $(X_{1},Y_{1})$ the random variable such that $(X_{1},Y_{1})\equiv(X,Y|E)$. Since
$P(E)$ $=P(|(X,Y)|\geq\eta)=1-P\big{(}(X,Y)\in B((0,0),\eta)\big{)}$
$\geq 1-C\eta^{s_{1}+s_{2}}\geq 1-C\eta^{2s}>\frac{1}{2},$
Lemma A.2 shows that the random variables $X_{1}$, $Y_{1}$, $(X_{1},Y_{1})$ are $(s_{1},O(C\cdot\max\{1,c\}))$-Frostman, $(s_{2},O(C\cdot\max\{1,c\}))$-Frostman and $(s_{1}+s_{2},O(C)))$-Frostman, respectively. Combining Lemma 2.24(i), one sees that
$\begin{split}H_{n}(X_{1})+H_{n}(Y_{1})&\geq n(s_{1}+s_{2})-2\log C+O(\log_{+}c)+O(1),\\ H_{n}(X_{1},Y_{1})&\geq n(s_{1}+s_{2})-\log C+O(1).\end{split}$ (42)
Since $s_{1}+s_{2}\geq 2s$, the right-hand sides of both inequalities in (42) are at least $2ns+O(\log_{+}C)+O(\log_{+}c)+O(1)$.
On the other hand, We can apply Lemma A.2, then combine (41) to get that
$\begin{split}&\frac{H_{n}(X_{1})+H(Y_{1})}{2}\leq\frac{H_{n}(X)+H_{n}(Y)}{2P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi),\\ &H_{n}(X_{1}+Y_{1})\leq\frac{H_{n}(X+Y)}{P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi),\\ &H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})\leq\frac{H_{n}(\alpha X^{2}+\beta Y^{2})}{P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi).\end{split}$ (43)
@@
$\begin{split}&H_{n}(X_{1})+H_{n}(Y_{1})-H_{n}(X_{1},Y_{1})=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1),\\ &H_{n}(X_{1}+Y_{1})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1),\\ &H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1).\end{split}$
Now we write $\log K:=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1)$. Applying Theorem 3.1 with the random variable $(X_{1},Y_{1})$, we obtain the random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$, where $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$, and
$H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(X_{1})-\log K,\quad H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(Y_{1})-\log K.\\$ (44)